OpenStudy (anonymous):

h = -16t2 + 36t + 4.

OpenStudy (anonymous):

what are you suppose to do in the equation?

OpenStudy (anonymous):

simplify?

OpenStudy (anonymous):

idk this is the full question If a football player passes a football from 4 feet off the ground with an initial velocity of 36 feet per second, how long will it take the football to hit the ground? Use the equation h = -16t2 + 36t + 4.

OpenStudy (anonymous):

That's a Quadratic Equation.

OpenStudy (anonymous):

\[\Large \frac{9+\sqrt{97}}{8} \ \text{or} \ \frac{9-\sqrt{97}}{8}\]

OpenStudy (anonymous):

\[\Large \frac{9+\sqrt{97}}{8} \approx 2.356107225\] \[\Large \frac{9-\sqrt{97}}{8} \approx -0.1061072252\]

OpenStudy (anonymous):

Then toss out the negative answer since it doesn't make any sense. So the ball will hit the ground at about 2.356 seconds

OpenStudy (anonymous):

yes that is one of the answer choices. good job! thank you